Grade 12th passPhysical Chemistry

How much water should be added to 200cc of seminormal solution of NaOH to make it exactly decinormal ??

Gaurav
8 Years agoGrade 12th pass

1 Answer

Khimraj
8 Years ago
200 cc = 100ml
seminormal = N/2
so (equivalent of NaOH)/0.2 = ½
so equivalent of NaOH = 0.1
let volume of water is X L
then
0.1/(0.2 + X) = 1/10
1 = 0.2 + x
s0 x = 0.8 L = 800ml = 800cc
Hope it clears.